# On the Hausdorff dimension and singularities of the monopolist’s free boundary curve

### 14 Mar 2026 v1

@@ Abstract.
The simplest genuinely multidimensional monopolist’s problem involves minimizing a linearly perturbed Dirichlet energy among nonnegative convex functions $u$ on an open domain $X\subset[0,\infty)^{2}$. The geometry of the region of strict convexity $\Omega\subset X$ for the unique minimizer $u$ is of central interest. A relatively closed portion $X_{1}^{0}\subset X$ of the domain is comprised of segments starting and ending on $\partial X$ along which $u$ is affine. For convex polygons and certain other domains $X\subset{\mathbf{R}}^{2}$, we build on results with Zhang [22] to show that outside $X_{1}^{0}\cup\{u=0\}$, the free boundary of $\Omega$ is a continuous curve of Hausdorff dimension one, and that $\Omega$ has density $1/2$ along it (and is $C^{\alpha}_{\mathrm{loc}}$ for all $0<\alpha<1$), except perhaps at a discrete set of singular points. We do this by showing that much of the free boundary solves an obstacle problem whose endogenous obstacle is $C^{2}$. For Rochet and Choné’s square $X=(a,a+1)^{2}$ with $a>0$, there is a point $x_{a}$ on the diagonal such that $X_{1}^{0}\cap\partial\Omega\subset\{x_{a}\}$, and the discrete singularities mentioned above can only accumulate at $x_{a}$ or at the two ends of the analytic arc $\{u=0\}\cap\partial\Omega$, plus the two limit points of $X\cap\partial\Omega$ on $\partial X$. Where the regularity of the endogenous obstacle can be improved to $C^{2,\mathrm{Dini}}$, the free boundary becomes locally $C^{\infty}$ outside a closed set whose relative interior is empty.
@@
Robert McCann’s work was supported in part by the Canada Research Chairs program CRC-2020-00289 and Natural Sciences and Engineering Research Council of Canada Discovery Grants RGPIN–2020–04162. Lucas O’Brien’s work was supported in part by a University of Toronto Excellence Award and by an internal fellowship from the Massachusetts Institute of Technology School of Science. Cale Rankin’s work was supported by grants DP220100067 & DE260100829 of the Australian Research Council. The authors are grateful to Shibing Chen and Kelvin Shuangjian Zhang for fruitful exchanges.
@@ 1. Introduction
of convex indirect utilities; here $\mathcal{H}^{k}$ denotes $k$-dimensional Hausdorff measure. The product selected turns out to depend continuously on consumer type $x\in X$ [34, 9], and to be given by $\nabla u(x)$, where $u$ is the unique maximizer of (1.1)–(1.2). We assume this continuity extends to $\overline{X}$ as in [9, 34]; see however [22, Remark 1.9]. Generalizations to other distributions of types [34], preferences [35, 10, 17] and transfers [32, 24], including regularity results [12, 23], have also been given e.g. by Carlier, Chen, Figalli, Kim, McCann, Noldeke, Rankin, Rochet, Samuelson, Zhang, and to auction design by Daskalakis, Deckelbaum and Tzamos [13]. Rochet–Choné showed the maximizer of (1.1)–(1.2) was characterized by the convex–ordering of the positive and negative parts of the variational derivative, conditioned on the action selected. A duality-based characterization [13, 25] was eventually parlayed into a characterization via Euler-Lagrange equations for a free boundary problem by two of us with Zhang [22]. However, none of these characterizations extend to the more general settings of [10, 17, 32, 24].
For convex domains $X\subset{\mathbf{R}}^{n}$, a deeper analysis of the maximizer of (1.1)–(1.2) performed by two of us with Zhang [22] revealed unexpected structures, only partly foreshadowed by numerics of Rochet–Choné [34], Ekeland–Moreno-Bromberg [14], Mirebeau [27], Boerma–Tsyvinsky–Zimin [5], Carlier–Dupuis–Rochet–Thanassoulis [11] and theory of McCann and Zhang [26, 25]. For $n=2$ the analysis of the free boundary separating the customization region — the largest open set $\Omega_{n}\subset X$ on which $u$ is strictly convex — from its complement was initiated in [22]; it will be continued here, still for $n=2$. Segments along which $u$ is affine that have both endpoints on $\partial X$ form a relatively closed subset of $X$. For domains satisfying a family of regularity assumptions (see §1.1(a)–(d)) — which include all convex polygons $X\subset[0,\infty)^{2}$ — we shall show that outside of this subset, the free boundary $X\cap\partial\Omega_{2}$ has Hausdorff dimension 1, and $\Omega_{2}$ has density $1/2$ along $X\cap\partial\Omega_{2}$, apart from a discrete set of singular points. This is achieved by building on our results with Zhang [22] and combining them with techniques pioneered for the obstacle problem with $C^{2,\alpha}$ obstacles by Caffarelli [8, 6], and extended to lower regularity obstacles by Blank [4]. Since our obstacle arises endogenously from $u$ and was not known to be better than $C^{1,1}$ previously, we use [22]*Lemma 7.3 and its proof to show the obstacle is $C^{2}$ and control the singular points on the free boundary before analyzing the regular points (where $\Omega_{2}$ has density $1/2$) by exploiting developments in Blank [4].
@@ 1.1. Statement of main results
Assumptions (a)–(d) are satisfied on convex polygons $X\subset[0,\infty)^{2}$ [22, Theorems 1.1, 4.1 and Remarks 1.9, 8.4], and potentially also on a much wider class of domains. Since a reflection and strict maximum principle argument [22, Lemma 3.2] for Poisson’s equation implies $(\tilde{x}\cap\partial\Omega_{2})\leq 1$ for each $x\in X_{1}^{0}$ in all the examples that we are aware of, it is natural to conjecture (e) holds quite generally. We have not verified this conjecture unless $X=(a,a+1)^{2}$, in which case $(X_{1}^{0}\cap\partial\Omega_{2})\leq 1$ by Remark 1.3.
@@
Let $u$ maximize the monopolist’s profits (1.1)–(1.2) on an open bounded convex $X\subset\mathbf{R}^{2}$ satisfying (a)-(c) and define the regular $\mathcal{R}$ and singular $\mathcal{S}$ parts of the tame free boundary $\mathcal{T}:=X_{1}^{+}\cap\partial\Omega_{2}$ by (1.3)–(1.7). Then $u\in C^{2}(X_{1}^{+})$, $\mathcal{T}$ is relatively open in $\partial\Omega_{2}$, and
@@
Proof of (1). To begin, note the decomposition of $\overline{X}$ as the union $\overline{X}={X_{0}\cup X_{1}^{-}\cup X_{1}^{0}\cup X_{1}^{+}\cup X_{2}\cup\partial X}$ yields a partition of $\partial\Omega_{2}\setminus X^{0}_{1}$ into components $X_{0}\cap\partial\Omega_{2}$, $X_{1}^{\pm}\cap\partial\Omega_{2}$, and $\partial X\cap\partial\Omega_{2}$ since [22, Theorem 1.1] shows $X_{2}$ is open. We will prove each of these components has Hausdorff dimension $1$. First, $X_{1}^{+}\cap\partial\Omega_{2}$ has Hausdorff dimension equal to $1$ by Theorem 1.1(2). Next, the boundaries of the convex sets $X$ and $\Omega_{0}=\{u=0\}$ are each covered by Lipschitz curves. Finally assumption (d) implies $(\overline{X_{1}^{-}}\cap\partial\Omega_{2})\setminus(X_{0}\cup X^{0}_{1})$ is countable.
@@
First, the intersection of $\overline{\mathcal{S}^{\prime}}$ with $X_{1}^{+}$ is the discrete set quantified by Theorem 1.1(1). Next, $(\overline{\mathcal{S}^{\prime}}\cap\overline{X_{1}^{-}})\setminus(X_{0}\cup X^{0}_{1})$ is closed and countable by hypothesis (d). We claim the set $FB\cap\overline{X_{1}^{-}}\cap\partial\Omega_{0}$ is empty. For a contradiction take $x\in FB\cap\overline{X_{1}^{-}}\cap\partial\Omega_{0}$, then $x$ is a limit of stray rays with length bounded below and so lies in a line segment contained in the boundary of $\Omega_{0}$. Since at most one point in this line segment, namely the endpoint, lies in $\overline{X_{1}^{+}}$ and the stray rays have measure 0, we see provided $x$ is not the endpoint of the line segment in $\partial\Omega_{0}$, $B_{\varepsilon}(x)\setminus(\Omega_{0}\cup\Omega_{2})$ has measure 0 for some $\varepsilon>0$. Thus, a maximum principle and reflection argument, as in [22, Lemma 3.2], yields a contradiction.
The remaining subset of $\overline{\mathcal{S}^{\prime}}$ to consider is elements in $U:=X\cap\partial\Omega_{0}\setminus\overline{X_{1}^{0}\cup X_{1}^{-}}$. We first establish that $\Omega_{0}$, which is empty or satisfies $\Omega_{0}=\{u=0\}$ by [22, Theorem 1.1], has $\Omega_{0}\cap\partial X$ connected. This follows from considerations laid out in Rochet and Choné [34, Proposition 1]: reflection symmetry across the axes and concavity of the principals’ maximization problem implies $Du\in[0,\infty)^{2}$, hence $X\cap\partial\Omega_{0}$ is a graph over the antidiagonal of a concave curve with Lipschitz constant at most $1$. Next, $x\cdot\bf n\leq 0$ along $\Omega_{0}\cap\partial X$ by [22, Proposition 2.3], which implies the curve $\Omega_{0}\cap\partial X$ intersects only the portion of $\partial X$ which is parameterized as a convex curve over the antidiagonal. Since the difference of a concave and convex function which does not vanish on an interval can have at most two zeros, it follows that $\Omega_{0}\cap\partial X$ is connected.
@@
The relative interiors of these connected components are analytic curves disjoint from the singular set, by [22, Remark 1.2]. We’ve established each set on the right-hand side of (1.8) has a countable intersection with $\overline{\mathcal{S}^{\prime}}$, thereby completing the proof of (2).
@@
Assertions (4)-(5) are straightforward consequences of (1)-(3) and the additional hypothesis (e). ∎
@@
In the case of the square $X=(a,a+1)^{2}$ with $a>0$, the preceding corollary yields $\mathrm{dim}_{\mathcal{H}}(\partial\Omega_{2})\leq 1$ and $\overline{\mathcal{S}^{\prime}}$ at most countable since $(X_{1}^{0}\cap\partial\Omega_{2})\leq 1$ and $(\partial X\cap\overline{X\cap\partial\Omega_{2}})\leq 2$ by [22, Theorem 1.5 and Lemma 8.7]. The square’s geometry models buyers whose two characteristics $x=(x^{1},x^{2})$ are distributed independently and uniformly over two intervals; the economic relevance of this has made it a prominent but challenging test case for numerics and theory [14, 27, 5, 11, 22, 25, 26].
@@ 3.3. Singular points are discrete
The main theorem of this section shows that set of singular points of the monopolist’s free boundary is a discrete subset of the tame free boundary, i.e. there is no tame free boundary point which is the limit of singular tame free boundary points. Before proceeding, we need to show that every singular point is a local maximum for $R$ — and not just the accumulation point of such maxima as implied by [22, Theorem 7.8].
@@ 4.2. Dini condition gives partial regularity at tame free boundary points
Proposition 4.6 (Conditional partial regularity of the tame free boundary).
Take $u_{1}$ to be defined as in Lemma 2.3, and assume that $\Delta u_{1}$ is Dini continuous. Then locally $\mathcal{T}$ is a $C^{1}$ curve except at the relatively discrete set $\mathcal{S}$, and is $C^{\infty}$ outside of a closed, nowhere dense subset of $\mathcal{T}$.
@@
Since Dini continuity of $\Delta u_{1}$ implies that $F(v)$ is $C^{1}$ in a neighbourhood of any regular point [4]*Theorem 0.1, it follows that $\mathcal{T}$ is a $C^{1}$ curve except at the relatively discrete subset $\mathcal{S}\subset\mathcal{T}$ of singular points. Now, consider any connected component $A$ of the set $\mathcal{R}$ of regular points, then $A$ is a $C^{1}$ curve by [4]*Theorem 0.1. Choose the tangent vector $\tau:A\to\mathbf{S}^{1}$ continuously. Since the ray direction $\xi$ is Lipschitz by [22]*Lemma 6.5, the map $x\mapsto|\xi(x)\cdot\tau(x)|$ is continuous, so the set
@@
So, since $\xi$ is Lipschitz, we may apply the Picard-Lindelöf theorem to conclude that $\eta(t)=f(t)$. Therefore, for any $x_{0}\in W$, we have that $W$ is contained entirely in the ray $\widetilde{x_{0}}$, contradicting the continuity of $R$. ∎

### 29 Mar 2026 v2

@@ Abstract.
The simplest genuinely multidimensional monopolist’s problem involves minimizing a linearly perturbed Dirichlet energy among nonnegative convex functions $u$ on an open domain $X\subset[0,\infty)^{2}$. The geometry of the region of strict convexity $\Omega\subset X$ for the unique minimizer $u$ is of central interest. A relatively closed portion $X_{1}^{0}\subset X$ of the domain is comprised of line segments starting and ending on $\partial X$ along which $u$ is affine. For convex polygons and potentially all domains $X\subset{\mathbf{R}}^{2}$, we build on results with Zhang to show that outside $X_{1}^{0}\cup\{u=0\}$, the free boundary of $\Omega$ is a continuous curve of Hausdorff dimension one, and that $\Omega$ has density $1/2$ along it (and is $C^{\alpha}_{\mathrm{loc}}$ for all $0<\alpha<1$), except perhaps at a discrete set of singular points. We do this by showing that much of the free boundary solves an obstacle problem whose endogenous obstacle is $C^{2}$.
From a slightly stronger conclusion, we deduce the free boundary becomes locally $C^{\infty}$ outside a closed set whose relative interior is empty. In response to the circulation of the present manuscript, we received a concurrent but independent work of Chen, Figalli and Zhang who verify a strengthening sufficient for this partial regularity result; (they show in particular that $\alpha=1$ and the discrete set mentioned above is empty).
@@
Robert McCann’s work was supported in part by the Canada Research Chairs program CRC-2020-00289 and Natural Sciences and Engineering Research Council of Canada Discovery Grants RGPIN–2020–04162. Lucas O’Brien’s work was supported in part by a University of Toronto Excellence Award and by an internal fellowship from the Massachusetts Institute of Technology School of Science. Cale Rankin’s work was supported by grants DP220100067 & DE260100829 of the Australian Research Council. The authors are grateful to Shibing Chen, Alessio Figalli, Yi Ru-Ya Zhang, and Kelvin Shuangjian Zhang for fruitful exchanges.
@@ 1. Introduction
of convex indirect utilities; here $\mathcal{H}^{k}$ denotes $k$-dimensional Hausdorff measure. If $X$ is convex, the product selected turns out to depend continuously on consumer type $x\in X$ [35, 9], and to be given by $\nabla u(x)$, where $u$ is the unique maximizer of (1.1)–(1.2). This continuity also extends to $\overline{X}$ [12]. Generalizations to other distributions of types [35], preferences [36, 10, 18] and transfers [33, 25], including regularity results [13, 24], have also been given e.g. by Carlier, Chen, Figalli, Kim, McCann, Noldeke, Rankin, Rochet, Samuelson, Zhang, and to auction design by Daskalakis, Deckelbaum and Tzamos [14]. Rochet–Choné showed the maximizer of (1.1)–(1.2) was characterized by the convex–ordering of the positive and negative parts of the variational derivative, conditioned on the action selected. A duality-based characterization [14, 26] was eventually parlayed into a characterization via Euler-Lagrange equations for a free boundary problem by two of us with Zhang [23]. However, none of these characterizations extend to the more general settings of [10, 18, 33, 25].
For convex domains $X\subset{\mathbf{R}}^{n}$, a deeper analysis of the maximizer of (1.1)–(1.2) performed by two of us with Zhang [23] revealed unexpected structures, only partly foreshadowed by numerics of Rochet–Choné [35], Ekeland–Moreno-Bromberg [15], Mirebeau [28], Boerma–Tsyvinsky–Zimin [5], Carlier–Dupuis–Rochet–Thanassoulis [11] and theory of McCann and Zhang [27, 26]. For $n=2$ the analysis of the free boundary separating the customization region — the largest open set $\Omega_{n}\subset X$ on which $u$ is strictly convex — from its complement was initiated in [23]; it will be continued here, still for $n=2$. Line segments along which $u$ is affine that have both endpoints on $\partial X$ form a relatively closed subset of $X$. For domains satisfying a family of regularity assumptions (see §1.1(a)–(d)) — which include all convex polygons $X\subset[0,\infty)^{2}$ — we shall show that outside of this subset, the free boundary $X\cap\partial\Omega_{2}$ has Hausdorff dimension 1, and $\Omega_{2}$ has density $1/2$ along $X\cap\partial\Omega_{2}$, apart from a discrete set of singular points. This is achieved by building on our results with Zhang [23] and combining them with techniques pioneered for the obstacle problem with $C^{2,\alpha}$ obstacles by Caffarelli [8, 6], and extended to lower regularity obstacles by Blank [4]. Since our obstacle arises endogenously from $u$ and was not known to be better than $C^{1,1}$ previously, we use [23]*Lemma 7.3 and its proof to show the obstacle is $C^{2}$ and control the singular points on the free boundary before analyzing the regular points (where $\Omega_{2}$ has density $1/2$) by exploiting developments in Blank [4].
@@ 1.1. Statement of main results
Assumptions (a)–(c) are satisfied on all convex domains [12, Theorem 1.1] [23, Theorem 4.1], while (d) holds at least on convex polygons $X\subset[0,\infty)^{2}$ [23, Remark 8.4]. Since a reflection and strict maximum principle argument [23, Lemma 3.2] for Poisson’s equation implies $(\tilde{x}\cap\partial\Omega_{2})\leq 1$ for each $x\in X_{1}^{0}$ in all the examples that we are aware of, it is natural to conjecture (e) holds quite generally. We have not verified this conjecture unless $X=(a,a+1)^{2}$, in which case $(X_{1}^{0}\cap\partial\Omega_{2})\leq 1$ by Remark 1.3.
@@
Let $u$ maximize the monopolist’s profits (1.1)–(1.2) on an open bounded convex $X\subset\mathbf{R}^{2}$ satisfying (a)–(c) and define the regular $\mathcal{R}$ and singular $\mathcal{S}$ parts of the tame free boundary $\mathcal{T}:=X_{1}^{+}\cap\partial\Omega_{2}$ by (1.3)–(1.7). Then $u\in C^{2}(X_{1}^{+})$, $\mathcal{T}$ is relatively open in $\partial\Omega_{2}$, and
@@
Proof of (1). To begin, note the decomposition of $\overline{X}$ as the union $\overline{X}={X_{0}\cup X_{1}^{-}\cup X_{1}^{0}\cup X_{1}^{+}\cup X_{2}\cup\partial X}$ yields a partition of $\partial\Omega_{2}\setminus X^{0}_{1}$ into components $X_{0}\cap\partial\Omega_{2}$, $X_{1}^{\pm}\cap\partial\Omega_{2}$, and $\partial X\cap\partial\Omega_{2}$ since [23, Theorem 1.1] shows $X_{2}$ is open. We will prove each of these components has Hausdorff dimension $1$. First, $X_{1}^{+}\cap\partial\Omega_{2}$ has Hausdorff dimension equal to $1$ by Theorem 1.1(2). Next, the boundaries of the convex sets $X$ and $\Omega_{0}=\{u=0\}$ are each covered by Lipschitz curves. Finally assumption (d) implies $(\overline{X_{1}^{-}}\cap\partial\Omega_{2})\setminus(X_{0}\cup X^{0}_{1})$ is countable.
@@
First, the intersection of $\overline{\mathcal{S}^{\prime}}$ with $X_{1}^{+}$ is the discrete set quantified by Theorem 1.1(1). Next, $(\overline{\mathcal{S}^{\prime}}\cap\overline{X_{1}^{-}})\setminus(X_{0}\cup X^{0}_{1})$ is closed and countable by hypothesis (d). We claim the set $FB\cap\overline{X_{1}^{-}}\cap\partial\Omega_{0}$ is empty. For a contradiction take $x\in FB\cap\overline{X_{1}^{-}}\cap\partial\Omega_{0}$, then $x$ is a limit of stray rays with length bounded below and so lies in a line segment contained in the boundary of $\Omega_{0}$. Since at most one point in this line segment, namely the endpoint, lies in $\overline{X_{1}^{+}}$ and the stray rays have measure 0, we see provided $x$ is not the endpoint of the line segment in $\partial\Omega_{0}$, $B_{\varepsilon}(x)\setminus(\Omega_{0}\cup\Omega_{2})$ has measure 0 for some $\varepsilon>0$. Thus, a maximum principle and reflection argument, as in [23, Lemma 3.2], yields a contradiction.
The remaining subset of $\overline{\mathcal{S}^{\prime}}$ to consider is elements in $U:=X\cap\partial\Omega_{0}\setminus\overline{X_{1}^{0}\cup X_{1}^{-}}$. We first establish that $\Omega_{0}$, which is empty or satisfies $\Omega_{0}=\{u=0\}$ by [23, Theorem 1.1], has $\Omega_{0}\cap\partial X$ connected. This follows from considerations laid out in Rochet and Choné [35, Proposition 1]: reflection symmetry across the axes and concavity of the principals’ maximization problem implies $Du\in[0,\infty)^{2}$, hence $X\cap\partial\Omega_{0}$ is a graph over the antidiagonal of a concave curve with Lipschitz constant at most $1$. Next, $x\cdot\bf n\leq 0$ along $\Omega_{0}\cap\partial X$ by [23, Proposition 2.3], which implies the curve $\Omega_{0}\cap\partial X$ intersects only the portion of $\partial X$ which is parameterized as a convex curve over the antidiagonal. Since the difference of a concave and convex function which does not vanish on an interval can have at most two zeros, it follows that $\Omega_{0}\cap\partial X$ is connected.
@@
The relative interiors of these connected components are analytic curves disjoint from the singular set, by [23, Remark 1.2]. We’ve established each set on the right-hand side of (1.8) has a countable intersection with $\overline{\mathcal{S}^{\prime}}$, thereby completing the proof of (2).
@@
Assertions (4)–(5) are straightforward consequences of (1)–(3) and the additional hypothesis (e). ∎
@@
In the case of the square $X=(a,a+1)^{2}$ with $a>0$, the preceding corollary yields $\mathrm{dim}_{\mathcal{H}}(\partial\Omega_{2})\leq 1$ and $\overline{\mathcal{S}^{\prime}}$ at most countable since $(X_{1}^{0}\cap\partial\Omega_{2})\leq 1$ and $(\partial X\cap\overline{X\cap\partial\Omega_{2}})\leq 2$ by [23, Theorem 1.5 and Lemma 8.7]. The square’s geometry models buyers whose two characteristics $x=(x^{1},x^{2})$ are distributed independently and uniformly over two intervals; the economic relevance of this has made it a prominent but challenging test case for numerics and theory [15, 28, 5, 11, 23, 26, 27].
The following remark improves the conditional result of Theorem 1.1(4) by showing $C^{\infty}$-regularity outside a topologically small set unconditionally.
Remark 1.4 (Note added in revision).
In response to the circulation of the present manuscript, we received a concurrent but independent work of Chen, Figalli and Zhang [12] that, in addition to verifying the boundary regularity required for (c), improves on Theorem 1.1(1),(3) for $\partial X\in C^{1,1}$ showing that $\mathcal{S}=\emptyset$ and $\alpha=1$. Under (a)–(b), this improved regularity combines with our proof of Proposition 4.6 to yield a better conclusion than Theorem 1.1(4): namely, unconditional smoothness of the tame free boundary $\mathcal{T}$ outside of a closed subset whose interior relative to $\mathcal{T}$ is empty (independent of whether or not $\Delta u$ is Dini continuous throughout $X_{1}^{+}$).
@@ 3.3. Singular points are discrete
The main theorem of this section shows that set of singular points of the monopolist’s free boundary is a discrete subset of the tame free boundary, i.e. there is no tame free boundary point which is the limit of singular tame free boundary points. Before proceeding, we need to show that every singular point is a local maximum for $R$ — and not just the accumulation point of such maxima as implied by [23, Theorem 7.8].
@@ 4.2. Dini condition gives partial regularity at tame free boundary points
Proposition 4.6 (Partial regularity of the tame free boundary).
Take $u_{1}$ to be defined as in Lemma 2.3.
  1. (1)
    If a relatively open subset $A\subseteq\mathcal{T}$ is a $C^{1}$ curve, then $A$ is $C^{\infty}$ outside of a closed, nowhere dense subset of $\mathcal{T}$.
  2. (2)
    If instead $\Delta u_{1}$ is Dini continuous, then $\mathcal{T}$ is a $C^{1}$ curve except at the relatively discrete set $\mathcal{S}$, and is $C^{\infty}$ outside of a closed, nowhere dense subset of $\mathcal{T}$.
@@
(1) Suppose some relatively open $A\subseteq\mathcal{T}$ is a $C^{1}$ curve. Choose the tangent vector $\tau:A\to\mathbf{S}^{1}$ continuously. Since the ray direction $\xi$ is Lipschitz by [23]*Lemma 6.5, the map $x\mapsto|\xi(x)\cdot\tau(x)|$ is continuous, so the set
@@
So, since $\xi$ is Lipschitz, we may apply the Picard-Lindelöf theorem to conclude that $\eta(t)=f(t)$. Therefore, for any $x_{0}\in W$, we have that $W$ is contained entirely in the ray $\widetilde{x_{0}}$, contradicting the continuity of $R$.
(2) Now, assume $\Delta u_{1}$ is Dini continuous. Since Dini continuity of $\Delta u_{1}$ implies that $F(v)$ is $C^{1}$ in a neighbourhood of any regular point [4]*Theorem 0.1, it follows that $\mathcal{T}$ is a $C^{1}$ curve except at the relatively discrete subset $\mathcal{S}\subset\mathcal{T}$ of singular points. We have therefore shown that every connected component $A$ of $\mathcal{T}\setminus\mathcal{S}$ is $C^{\infty}$ outside of a closed, nowhere dense subset of $\mathcal{T}$, so we conclude that $\mathcal{T}$ is $C^{\infty}$ except at a closed, nowhere dense subset. ∎
The result of Proposition 4.6(1) becomes unconditional in view of Remark 1.4.